OpenStudy (anonymous):

7. Kepler’s Third Law of Orbital Motion states that you can approximate the period P (in Earth years) it takes a planet to complete one orbit of the sun using the function , where d is the distance (in astronomical units, AU) from the planet to the sun. How many Earth years would it take for a planet that is 6.76 AU from the sun? (1 point)15.23 17.58 154.46 3.58

OpenStudy (anonymous):

This one was hard. \[P=6.76^{\frac{ 3 }{ 2 }}\]

OpenStudy (anonymous):

whats 3/2?

OpenStudy (anonymous):

1 1/2

OpenStudy (anonymous):

good, so 1.5. now it looks like this \[P=6.76^{1.5}\]

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

then do 6.76^1.5 and you'll get the answer

OpenStudy (anonymous):

MULTIPLY 6.76 1.50

OpenStudy (anonymous):

DO I MULTIPLY THEM

OpenStudy (anonymous):

no 1.5 is the exponent

OpenStudy (anonymous):

THEN WHAT HOW DO I GET THE ANSWER

OpenStudy (anonymous):

i told you do \[6.76^{1.5}\]

OpenStudy (anonymous):

3.58

OpenStudy (anonymous):

no

OpenStudy (anonymous):

17.58

OpenStudy (anonymous):

yes, good job

OpenStudy (anonymous):

5. Simplify . (1 point) 9 81 5. Simplify . (1 point) 9 81

OpenStudy (anonymous):

9 1/3 81 1/3

OpenStudy (anonymous):

3.d 4.b 5.a

OpenStudy (anonymous):

no, none of those are right,

OpenStudy (anonymous):

3. p=2l+2w \[p= 2(7\sqrt{13})+2(\sqrt{13})\] \[p=14\sqrt{13}+2\sqrt{13}\] \[p=16\sqrt{13}\]

OpenStudy (anonymous):

16 square root 13 so its b

OpenStudy (anonymous):

yepp

OpenStudy (anonymous):

i dont get # 4

OpenStudy (anonymous):

you have to factor oout the numbers like we did with the cube root earlier but this time its only square root so you only need groups of 2 of the same number

OpenStudy (anonymous):

like for 20 it'd be 20 / \ 5 4 / \ 2 2

OpenStudy (anonymous):

2

OpenStudy (anonymous):

okay so then you put the 2 outside the radical and leave the five under it so it looks like this:\[2\sqrt{5}\]

OpenStudy (anonymous):

then do the same thing with the 45 45 / \ 5 9 / \ 3 3

OpenStudy (hungrylamp):

I just finished this test, scored 100 1. B: -4/7 and 4/7 2. C: 6|g^3| 3. B: 16(sqrt)13 units 4. A: 4(sqrt)5 5. B: 9 6. C: 8x ^-15/7 7. B: 17.58 8. A: x^3/5 For the last two, I have not been graded yet so use these at your own risk 9. -3 / (7 - 2(sqrt)10) (2 + 5) / (2 - 5) = (2 + 5) * (sr2 - sr5) / (sr2 - sr5) * (sr2 - sr5) = (sr2 - sr5) / (sr2 - sr5) = (2 - 5) / (sr2 - 2sr10+sr5) = -3 / (2 - 2(s)10 + 5) = -3 / (7 - 2(sqrt)10) 10. a^2/4